# The multimessenger view of Pulsar Timing Array black holes with the Horizon-AGN simulation

### v1

@@ Appendix C EM properties of MBHBs
In complement to Figure 14, where we show the electromagnetic properties of CW candidates, we report in Figure 15 an analogous figure for the MBHBs dominating the background. We model and analyze only the MBH binaries that are in the Horizon-AGN lightcone (Laigle et al. 2019). The sources are overall fainter than CW candidates, because located at higher redshift and characterized by lighter MBHs and galaxies (cf. Figure 9), but the results are in broad agreement, in the sense that generally the host galaxies outshine the MBHB emission at optical/near-IR wavelengths (cf. Figure 14, left) while for radiatively efficient accretors the MBHB emission is comparable or larger than the host galaxy’s in radio and X-rays. Bolometric luminosities tend to remain also in this case below quasars’ typical luminosities.
In Figure 16 we show the ratio of AGN and galaxy brightness for CW candidates. Results are qualitatively similar for binaries dominating the background. Red filled dots assume radiatively efficient accretion at all Eddington ratios, while empty red dots show a correction for radiatively inefficient emission for $f_{\rm Edd}<10^{-3}$. In optical and near-IR bands (B, V, and K) the host galaxies are always brighter than the MBHB-powered AGN for radiatively inefficient MBHs, and only a few of the efficient accretors have magnitudes comparable to or in a very few cases brighter than, their hosts with a preference for longer wavelengths (e.g., K), similarly to Veronesi et al. (2025). In radio and X-rays all efficient accretors ($f_{\rm Edd}>10^{-3}$) are brighter than the host, while the inefficient accretors are always out-shined by the stellar emission from the host galaxy in X-rays when including a correction for radiative efficiency. In radio, the fraction of inefficient accretors brighter than the hosts is 61%.

### v2

@@ Appendix C EM properties of MBHBs
In complement to Fig. 14, where we show the electromagnetic properties of CW candidates, we report in Fig. 15 an analogous figure for the MBHBs dominating the background. We model and analyze only the MBH binaries that are in the Horizon-AGN lightcone (Laigle et al. 2019). The sources are overall fainter than CW candidates, because they are located at higher redshift and characterized by lighter MBHs and galaxies (cf. Fig. 9), but the results are in broad agreement, in the sense that generally the host galaxies outshine the MBHB emission at optical/NIR wavelengths (cf. Fig. 14, left) while for radiatively efficient accretors the MBHB emission is comparable to or larger than the host galaxy’s in radio and X-rays. Bolometric luminosities tend to remain also in this case below quasars’ typical luminosities.
In Fig. 16 we show the ratio of AGN and galaxy brightness for CW candidates. Results are qualitatively similar for binaries dominating the background. Red filled dots assume radiatively efficient accretion at all Eddington ratios, while empty red dots show a correction for radiatively inefficient emission for $f_{\rm Edd}<10^{-3}$. In the optical and NIR bands (B, V, and K), the host galaxies are always brighter than MBHB-powered AGNs for radiatively inefficient MBHs, and only a few of the efficient accretors have magnitudes comparable to or, in a very few cases, brighter than, their hosts with a preference for longer wavelengths (e.g., K), similarly to Veronesi et al. (2025). In radio and X-rays all efficient accretors ($f_{\rm Edd}>10^{-3}$) are brighter than the host, while the inefficient accretors are always outshined by the stellar emission from the host galaxy in X-rays when including a correction for radiative efficiency. In radio, the fraction of inefficient accretors brighter than the hosts is 61%.